An investment had a nominal return of 10.3 percent last year. If the real return on the investment was only 5.7 percent, what was the inflation rate for the year?
Answer #1
(1+ Nominal Return ) = (1+ real return) (1+ Inflation rate)
(1+0.103) =(1+0.057)(1+ Inflation rate)
1.103 =1.057 * (1+ Inflation rate)
1.103/1.057 = 1+ Inflation rate
1.0435 = 1+ inflation rate
Inflation rate = 0.0435 i.e. 4.35%